OpenStudy (anonymous):

Use the discriminant to determine the number and type of solutions for the following equation. \[24x^2 + 13 = 0\] A. Zero real solutions B. One rational solution C. Two rational solutions D. Two irrational solutions

OpenStudy (anonymous):

I have no idea what to do. >_< @beenlightened

OpenStudy (anonymous):

ok do you know the method of PEMDAS?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

jk give me one moment to solve and explain

OpenStudy (solomonzelman):

For any \(\large\color{black}{ \displaystyle \color{blue}{a}{\rm x}^2+\color{blue}{c} =0 }\) \(\large\color{black}{ \displaystyle \color{blue}{a}{\rm x}^2=-\color{blue}{c} }\) \(\large\color{black}{ \displaystyle {\rm x}^2=\frac{-\color{blue}{c}}{\color{blue}{a}} }\) and that would be an imaginary solution.

OpenStudy (anonymous):

Thank you for your solution, but I had already done that much and I would like to ask, is the answer B.?? @SolomonZelman

OpenStudy (solomonzelman):

-c/a is a negative number correct?

OpenStudy (anonymous):

Correct.

OpenStudy (solomonzelman):

can you take a square root of a negative number (and get a real solution with this)?

OpenStudy (anonymous):

I don't believe so.

OpenStudy (anonymous):

The answer must be A. then.

OpenStudy (solomonzelman):

yes, correct, a square root of a negative number will not be equal to any real number

OpenStudy (solomonzelman):

yes, the answer is A.

OpenStudy (anonymous):

Got it. Thank you for your explanation. :)

OpenStudy (solomonzelman):

anytime:) and by the way the actual solution to this would be: \(\large\color{black}{ \displaystyle 24{\rm x}^2+13=0 }\) \(\large\color{black}{ \displaystyle 24{\rm x}^2=-13 }\) \(\large\color{black}{ \displaystyle {\rm x}^2=-\frac{13}{24} }\) \(\large\color{black}{ \displaystyle {\rm x}=\pm\sqrt{-\frac{13}{24}} }\) \(\large\color{black}{ \displaystyle {\rm x}=\pm\sqrt{-1}\times \sqrt{\frac{13}{24}} }\) \(\large\color{black}{ \displaystyle {\rm x}=\pm i\times \sqrt{\frac{13}{24}} }\) \(\large\color{black}{ \displaystyle {\rm x}=\pm i \sqrt{\frac{13}{24}} }\)

OpenStudy (anonymous):

Thank you for the solution. :)

OpenStudy (solomonzelman):

yes, that is an imaginary (not real) solution.

OpenStudy (solomonzelman):

you are welcome!